OpenStudy (anonymous):

In a group of 1200 people, 468 have brown eyes, 252 have blue eyes, 132 have green eyes, and 348 have black eyes. What is the probability that a person chosen at random from this group has blue eyes?

OpenStudy (paki):

see... The sample space here is 1200... lets call n(a) = number of people with blue eyes hence n(a) = 252.. Therefore p(a) = n(a)/sample space which is 252/1200... solve which is 0.2...

OpenStudy (anonymous):

subtract the ones you arent looking for and divide the remaining number by the total... from there you can solve for decimal percent or simplify the fraction.

OpenStudy (paki):

252/1200 = 21/100 = 0.21 = 21% @TaylorDyals

OpenStudy (anonymous):

yes exactly what paki did